# Robust variational neural posterior estimation for simulation-based inference

### v1

@@ 4.5 Task D: Spectra Task - Real Gaia BP/RP Data
The third data release of the European Space Agency’s Gaia telescope (Gaia Collaboration et al., 2016) contain over 220 million flux-calibrated, low-resolution, optical stellar spectra. These spectra are measured by two instruments, the “Blue Photometer” (BP, 330-680 nm coverage in wavelength) and the “Red Photometer” (RP, 640-1050 nm). The processed and calibrated (De Angeli, F. et al. 2023; Montegriffo, P. et al. 2023) BP/RP (XP) spectra from Gaia DR3 are low-resolution, contaminated spectra which have multiple difficult systematics to overcome (Huang et al., 2024). However, the XP spectra are expected to contain significant information about different stellar parameters (Witten et al., 2022) and robust posterior inference using stellar evolution simulators conditional on the Gaia XP spectra would provide an efficient method for understanding the Milky Way.
@@
Simulator We use the MIST Choi et al. (2016) stellar evolution models to generate stellar parameters compatible with high Galactic latitudes and map each of the effective temperature, log-surface gravity, and metallicity to a medium-resolution synthetic Castelli & Kurucz (2004) model spectral energy distribution. This defines the simulator relating $\bm{\theta}=(T_{\rm eff},\log g,\rm{[Fe/H]})$ to $\bm{x}_{\rm sim}$. We cut the spectra in their native resolution between 330-1050 nm to define a $301$ dimensional vector which overlaps with the Gaia XP spectra wavelength range but at a significantly different resolution.
@@
Embedding We pre-train an NSE estimator using the same method as described in the pendulum synthetic task.
@@
Misspecification: We can view the problem of inferring stellar parameters using real Gaia XP spectra and synthetic stellar evolution models as a misspecification problem. The Gaia XP spectra are expected to have two significant differences from the synthetic simulation model. Firstly, the XP spectra are lower resolution than the synthetic spectra and, therefore, contain less information than high-resolution spectra from the simulator. Secondly, there are inherent systematic errors in the processing of the XP spectra that cause a simulation gap even if we know the spectroscopic parameters using traditional high-resolution methods. For the real dataset, we target high Galactic latitudes due to the minimal impact of photometric extinction in these regions and a relatively homogenous population of main-sequence stars (O’Callaghan et al., 2024). We select all stars with absolute Galactic latitude $|b|>80^{\circ}$ that have valid LAMOST (Wang, 2022) spectroscopically determined stellar parameters. These spectroscopically determined stellar parameters will act as ground truth for our experiment, but we should note that they have their own errors arising from the spectroscopic determination. Furthermore, we select all Gaia recommended quality cuts and choose stars with confident distance estimates between $300$ and $700$ pc. This leaves us with a dataset of size $N_{\rm obs}=1053$.
Misspecification Summary. For those not familiar with stellar astronomy, the observed dataset is generated from real data with ground truth values that were selected so that measurement error and external factors have a minimal impact on the data, helping us better isolate the model misspecification.
Results. We display the results for the Spectra task in Figure 5 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2. Finally, we display the simulated observations together with observed points in Figure 10. In this task, the misspecification is geometrically very complex due to both the image of the neural statistic and the complexity of stellar evolution. Many of the points appear ID relative to the simulated points in this task. We naively applied the neural statistic, so that the neural statistic knows nothing about the structure of the Gaia XP spectra. We find that the neural statistic struggles to identify the metallicity parameter, most likely due to the neural statistic fitting for high-resolution features in the synthetic spectra.
Due to the complexity of the misspecification, NNPE struggles significantly. However, RVNP and RVNP-Global recover robust posterior inference for this task. In particular, RVNP-Global is very well calibrated after the $N_{\rm obs}=10$ point and recovers robust posterior inference across a broad range in parameter space. In this task, we displayed the tuned variant of RVNP-T and found that it struggled significantly to recover robust posterior inference. It is worth noting that we chose a poor prior for this problem, and we hypothesise that this is why RVNP-Global performs better than RVNP-Global, as the neural network covariance matrix accesses many unseen parts of parameter space.
@@
Figure 6: Results for the Spectra task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NMSE, lower values are better.
@@ 5 Related work
Misspecification in SBI Model misspecification is understood in likelihood-based methods (Davison 2003, pages 147–148); however, a systematic theory of likelihood-free methods is lacking. Misspecification in SBI has been studied in the context of ABC (Frazier et al. 2020; Bharti et al. 2022; Fujisawa et al. 2021), BSL (Frazier et al., 2024), generalized Bayesian inference (Dellaporta et al., 2022), and neural conditional density estimation (Ward et al. 2022; Kelly et al. 2024; Huang et al. 2023; Elsemüller et al. 2025; Mishra et al. 2025; Schmitt et al. 2024; Wehenkel et al. 2025). In this paper, we focused on neural-based methods, where empirically it has been shown that SBI struggles under model misspecification (Cannon et al. 2022; Schmitt et al. 2024). Robust posterior recovery under model misspecification is essential for the success of SBI, and different methods have emerged to mitigate against it. Kelly et al. (2025) identifies three main strategies currently used to account for model misspecification in SBI: robust summary statistics, generalised Bayesian inference, and error modelling and adjustment parameters. Most of the early solutions in robust SBI were intended for a single data set (Ward et al. 2022; Kelly et al. 2024; Huang et al. 2023).
@@
Robust amortised SBI Recently, robust amortised neural SBI has been addressed using optimal transport for domain shifts when a calibration set exists (Wehenkel et al., 2025), using unsupervised domain adaptation (Elsemüller et al., 2025), and consistency losses regularisation (Mishra et al., 2025). Moreover, Glöckler et al. (2023) proposes regularisation techniques to increase the robustness of the learnt posterior against adversarial attacks. Of these methods, only the consistency loss method targets both the likelihood and the posterior. The noisy neural posterior estimation from Ward et al. (2022) can also be viewed as an amortised SBI method, where a pre-defined error model is used during training to corrupt the simulations.
@@
Variational methods in SBI Variational methods have been used in multiple capacities to date. Wiqvist et al. (2021) introduced Sequential Neural Posterior and Likelihood Approximation, which proposes using variational inference (VI) to speed up the inference of likelihood-based methods, similar to the likelihood-based Bayesian approach to VI. Glöckler et al. (2022) introduces a framework that uses VI for simulation-based inference by using a pre-trained likelihood (or likelihood-ratio) and learn the posterior using VI, then refining the posterior using sampling importance resampling (Rubin, 1987). Nautiyal et al. (2024) introduces a generative modelling approach based on the variational inference framework and learns an encoder-decoder model in terms of latent variables. They introduce a latent variable that can account for complex structures and dependencies in the simulator model. Simons et al. (2022) propose a simulation-based inference algorithm that iteratively updates particles to more match the posterior in a variational likelihood-free gradient descent manner.
6 Discussion
In this paper, we introduced RVNP(-T), a robust amortised Bayesian inference method for simulation-based inference that jointly infers the simulation-to-reality gap and the amortised posterior using an importance-weighted autoencoder framework. This is the first case of using variational inference for robust simulation-based inference. Moreover, it is the first approach that does not rely on tuning hyperparameters of the loss function to recover robust posterior inference. We argue that this is an important step toward reliable posterior inference in amortised SBI. Previous work’s reliance on tuning parameters or misspecification priors to control the robustness of the posterior implies that there is an unknown parameter that controls the posterior inference in a way that may be nontrivial for real inference tasks. Furthermore, the error model is interpretable, overcoming the issues with unsupervised domain adaptation.
6.1 Neural statistic estimators
In this paper, we adopt a neural statistic estimator (NSE) before training the neural likelihood surrogate. There is an architectural limitation from the necessity of learning the likelihood proxy in RVNP because learning the NSE simultaneously with the likelihood is not a well-posed problem (Brehmer & Cranmer, 2020). However, we argue that in amortised SBI, this is not a significant problem. Usually, amortised SBI requires a fixed simulation budget up front, and in neural posterior estimation (NPE), the neural conditional density estimator is trained on those simulations. When training an NSE for an NPE task, usually the NSE is trained simultaneously with the posterior proxy. However, this objective function has no theoretical guarantees of the sufficiency of the statistic (Chen et al., 2021). In this paper, we chose to pre-train the NSE using the InfoMax objective (Hjelm et al., 2019) because of its theoretical guarantees of sufficiency.
We also considered hand-crafted statistics in this paper, as the robustness guarantees occur during the final variational inference step. This allows RVNP to include problems where an understanding of the error model is a significant part of the inference task. Such as in situations where the chosen summary statistics are in units known to the scientist, and they wish to fit for discrepancies between the model and the data.
6.2 Connection to domain adaptation
Unsupervised domain adaptation couples posterior learning to adapting the neural statistics. In this paper, we saw that for some tasks, learning a simple covariance matrix parameterised by a neural network on a subset of data does not necessarily generalise well to other areas of parameter space, even with strong inductive bias. Complex domain adaptation in a nonlinear fashion will give rise to infinite solutions that can account for the simulation-to-reality gap, and even in simple inductive bias error models throughout this paper, it highlights that the domain adaptation may cause serious issues if many of the points lie ID. A separate body of work should discuss error modelling using calibration to compare with the results of Wehenkel et al. (2025).
@@ 6.3 Prior misspecification
A type of misspecification not explicitly addressed in this paper is if the prior $p(\bm{\theta})$ is itself misspecified. RVNP should have the same limitations as likelihood-based prior misspecification, except with two crucial differences. Any prior chosen must be defined within the support of the prior used to train the likelihood. Furthermore, changing the prior distribution may make the likelihood proxy learn different parts of the parameter space suboptimally. In the Spectra task, we defined an incorrect prior relative to the observed data, which occurred unintentionally due to the selection effects of the real data. It is worth considering the pendulum task for this problem. The OOD points appear OOD under this misspecification because the effect of the misspecification makes the points appear they have a higher fundamental frequency. However, the information about the misspecification is only available to us because we assume that the prior is well-specified. We would lose the robustness of the inference if we assumed a wider prior on the fundamental frequency, causing the synthetic DGP to cover the observed data points.
6.4 Higher dimensional problems
Our method inherits the limitations of both neural likelihood estimation and amortised variational inference. It is highly desirable to use NSE in higher-dimensional problems to reduce the dimension of the data to help the computational capabilities of the normalising flow on the likelihood estimation. In general, RVNP should scale well to higher-dimensional problems, but may struggle to compete with domain adaptation methods for more complex, multimodal data. Future work will look at the limitations of using pre-training the NSE and learning the likelihood proxy on the embedded summary statistics.
Broader Impact Statement

### v2

@@ 4.5 Task D: Spectra Task - Real Gaia BP/RP Data
The third data release of the European Space Agency’s Gaia telescope (GAIA_MISSION) contain over 220 million flux-calibrated, low-resolution, optical stellar spectra. These spectra are measured by two instruments, the “Blue Photometer” (BP, 330-680 nm coverage in wavelength) and the “Red Photometer” (RP, 640-1050 nm). The processed and calibrated (bprpspec; external_bprp) BP/RP (XP) spectra from Gaia DR3 are low-resolution, contaminated spectra which have multiple difficult systematics to overcome (Huang_2024). However, the XP spectra are expected to contain significant information about different stellar parameters (2022MNRAS.516.3254W) and robust posterior inference using stellar evolution simulators conditional on the Gaia XP spectra would provide an efficient method for understanding the Milky Way.
@@
Simulator We use the MIST MIST stellar evolution models to generate stellar parameters compatible with high Galactic latitudes (a region where observable contaminants are low) and map each of the effective temperature, log-surface gravity, and metallicity to a medium-resolution synthetic 2003IAUS.210P.A20C model spectrum. This defines a stochastic simulator mapping $\bm{\theta}=(T_{\rm eff},\log g,\rm{[Fe/H]})$ to $\bm{x}_{\rm sim}$. We restrict the spectra in their native resolution between 330-1050 nm to define a $301$ dimensional vector which overlaps with the Gaia XP spectra wavelength, but at a significantly different resolution.
@@
Embedding We pre-train an NSE estimator using the same method as described in the pendulum task.
@@
Misspecification: We can view the problem of inferring stellar parameters using real Gaia XP spectra and synthetic stellar evolution models as a misspecification problem. The Gaia XP spectra are expected to have two significant differences from the synthetic simulation model. Firstly, the XP spectra are lower resolution than the synthetic spectra and, therefore, contain less information. Secondly, there are inherent systematic errors in the processing of the XP spectra and in the stellar evolution modelling that causes a simulation gap even if we knew the spectroscopic parameters exactly. For the real dataset, we target high Galactic latitudes due to the minimal impact of dust extinction and other contaminants (high_gal_me). We select all stars with absolute Galactic latitude $|b|>80^{\circ}$ that have valid LAMOST (LAMOSTDR8) spectroscopically determined stellar parameters. These spectroscopically determined stellar parameters will act as ground truth for our experiment, but we should note that they have their own errors arising from the spectroscopic determination. Furthermore, we select all Gaia recommended quality cuts and choose stars with confident distance estimates between $300$ and $700$ pc. This leaves us with a dataset of size $N_{\rm obs}=1053$.
Results. We show the results for the Spectra task in Figure 5 and provide an example posterior in the $N_{\rm obs}=1000$ case in Figure 2. We display the simulated observations together with the observed points in Figure 19. In Figure 11, we evaluate RVNP on a well-specified version of the Spectra task. In Figure 15, we plot the final two dimensions of the RVNP ($N_{\rm obs}=1000$) covariance matrix eigenvalues evaluated at different points of the training data.
In the spectra task, the misspecification is geometrically very complex. This arises from using a neural statistic embedding and from the complexity of stellar evolution models. Many of the points appear ID relative to the simulated points in this task. We naively applied the neural statistic, so that the neural statistic knows nothing about the structure of the Gaia XP spectra, including instrument response and measurement error. The RVNP and RVNP global algorithms can recover robust posterior inference beginning after the $N_{\rm obs}=10$ point. We find that both of these algorithms perform better than NNPE and NPE. However, RVNP-T collapses to the NPE posterior after tuning. NNPE struggles significantly, most likely due to the spike-and-slab error model being unable to account for such a complex misspecification. RVNP-G is very well calibrated after the $N_{\rm obs}=100$ point and recovers robust posterior inference across a broad range of parameter space. Furthermore, RVNP is reasonably well calibrated, particularly when considering the marginal calibration metric. NPE performs very poorly in the pendulum task in terms of calibration and log-probability, but its sampled points are close to the truth, reflecting an extreme overconfidence.
Investigating the covariance matrix of the error models helps us to understand why the global variant performs better. In certain points of parameter space where the covariance matrix has not observed data, the posterior collapses to the NPE estimate, and near these points the inferene is less reliable. The covariance matrix provides useful model criticism by considering the error as a function of $\bm{\theta}$. We find that we cannot recover the metallicity estimate reliably. This is not surprising because the resolution of the synthetic spectra is higher than that of the observed data, and the metallicity will appear as strong local features in the spectra. These features are not present in the Gaia BP/RP spectra. Domain adaptation approaches that can account for these difference may be desirable in this context.
@@
Figure 6: Results for the Spectra task. We conclude that RVNP and its variants can recover robust posterior inference in amortised simulation-based inference. The hue in the middle plots indicates the error bar on the NPE and NNPE algorithms. For $\alpha$ nearest to 0 is best, with positive values representing underconfidence and negative values representing overconfidence. For the log-probability, higher values are better. For NRMSE, lower values are better. (SIR) indicates sample-importance-resampling.
@@ 5 Related work
Misspecification in SBI Model misspecification is understood in likelihood-based methods (davison2003statistical, pages 147–148); however, a systematic theory of likelihood-free methods is lacking. Misspecification in SBI has been studied in the context of ABC (10.1111/rssb.12356; bharti2022approximate; fujisawa2021gamma), BSL (Frazier26092024), generalized Bayesian inference (dellaporta2022robust), and neural conditional density estimation (ward2022robust; kelly2024misspecificationrobust; huang2023learning; elsemüller2025doesunsuperviseddomainadaptation; mishra2025robust; schmitt2024detecting; wehenkel2025addressing). In this paper, we focused on neural-based methods, where empirically it has been shown that SBI struggles under model misspecification (cannon2022investigating; schmitt2024detecting). Robust posterior recovery under model misspecification is essential for the success of SBI, and different methods have emerged to mitigate against it. kelly2025simulation identifies three main strategies currently used to account for model misspecification in SBI: robust summary statistics, generalised Bayesian inference, and error modelling and adjustment parameters. Most of the early solutions in robust SBI were intended for a single data set (ward2022robust; kelly2024misspecificationrobust; huang2023learning).
@@
Robust amortised SBI Recently, robust amortised neural SBI has been addressed using optimal transport for domain shifts when a calibration set exists (wehenkel2025addressing), using unsupervised domain adaptation (elsemüller2025doesunsuperviseddomainadaptation), and consistency losses regularisation (mishra2025robust). Moreover, 2023arXiv230514984G proposes regularisation techniques to increase the robustness of the learnt posterior against adversarial attacks. Of these methods, only the consistency loss method targets both the likelihood and the posterior. The noisy neural posterior estimation from ward2022robust can also be viewed as an amortised SBI method, where a pre-defined error model is used during training to corrupt the simulations.
@@
Variational methods in SBI Variational methods have been used in multiple capacities to date. wiqvist2021sequential introduced Sequential Neural Posterior and Likelihood Approximation, which proposes using variational inference (VI) to speed up the inference of likelihood-based methods, similar to the likelihood-based Bayesian approach to VI. glockler2022variational introduces a framework that uses VI for simulation-based inference by using a pre-trained likelihood (or likelihood-ratio) and learn the posterior using VI, then refining the posterior using sampling importance resampling (Rubin1987SIR). nautiyal2024variational introduces a generative modelling approach based on the variational inference framework and learns an encoder-decoder model in terms of latent variables. They introduce a latent variable that can account for complex structures and dependencies in the simulator model. simons2022variational propose a simulation-based inference algorithm that iteratively updates particles to more match the posterior in a variational likelihood-free gradient descent manner. Moreover, dax2023neural use a trained NPE for importance-sampling in a likelihood-based manner.
6 Discussion and limitations
In this paper, we introduced RVNP(-T), a robust amortised Bayesian inference method for simulation-based inference that jointly infers the simulation-to-reality gap and the amortised posterior using an importance-weighted autoencoder framework. RVNP is the first example of using variational inference for robust simulation-based inference, and was inspired by robust SBI approaches for synthetic likelihood (frazier2021robust), neural posterior estimation (ward2022robust), and sequential neural likelihood estimation (kelly2024misspecificationrobust). We argue that RVNP’s novelty and its ability to recover robust amortised posterior inference without setting misspecification-related tuning parameters constitutes an important step toward reliable posterior inference in amortised SBI. Previous work’s reliance on setting a hyperparameter or misspecification prior to control the influence of the misspecification on the posterior inference is undesirable, and as more data are observed we will have a better understanding of the simulation-to-reality gap, which should inform the inference without strong assumptions on the degree of misspecification. Moreover, RVNP is reliable in the well-specified case.
Our work is not without limitations, and in what follows we discuss the scope of RVNP, how it relates to previous error modelling approaches, the use of neural statistic embeddings, and the implications of using variational inference to infer the posterior.
6.1 Error Modelling
Including an error model is an assumption about how to model the true data-generative process (DGP) under the assumption that observed data appears corrupted relative to the model simulations. In most real-world problems, the exact corruption is unknown and there are infinitely many distributions and assumptions that can bridge the simulation-to-reality gap. In this paper, we imposed a strong inductive bias on the distribution that bridges it, so that the error model accounts for the corruption by inflating the error on the synthetic DGP. This choice is a neural adaptation of the covariance inflation from frazier2021robust. The simulation-to-reality gap should inform our error model as we observed more data from the true DGP, particularly if the misspecification is a global phenomenon (Figure 1). By jointly inferring the posterior and the error model, we can allow the error model parameters to be inferred directly from the data without having tunable parameters that dictate the strength of the misspecification (for example, a shrinkage prior over the adjustment parameters (kelly2024misspecificationrobust) or the misspecification regularisation in unsupervised domain adaptation (huang2023learning)).
Similarly to the spike-and-slab error model introduced by ward2022robust, the error models introduced in this paper allows us to have a form of model criticism. We argue that without prior knowledge about the misspecification, inflating the error of the output of the simulator is the most sensible approach, as it is interpretable and allows for a level of trust in the inference procedure. The spike-and-slab model invokes a similar assumption: that the true corruption model is unknown and the misspecification encourages the posterior to forget certain summary statistics if they lie OOD. This relies on the misspecification appearing along individual axes in summary-statistic space, which is unrealistic in many situations because misspecification has no guarantee of appearing along given axes. Furthermore, the success of the spike-and-slab error model is subject to hyperparameter choices that dictate the strength of misspecification. We found that in the pendulum and the spectra task NNPE performed poorly for these reasons.
The parameter adjustment method of kelly2024misspecificationrobust can be thought of as an error model $p_{\bm{\xi}}(\bm{x}_{\rm obs}|\,\bm{x}_{\rm sim})=\delta(\bm{x}_{\rm obs}-\bm{x}_{\rm sim}-\bm{\xi})$ after integrating with respect to $\bm{x}_{\rm sim}$, where $\bm{\xi}$ are the linear adjustment parameters. This approach will move the observed point in summary-statistic space to a region of high probability with respect to the simulated samples. On a point-by-point basis, there are infinitely many solutions to this problem, and there is no guarantee of where an OOD point should be mapped back to in the original sample without extra information on the misspecification, such as a calibration set. Moreover, kelly2024misspecificationrobust state that the ability to recover robust posterior inference in their benchmark tasks depends on the hyperparameters that control the strength of the misspecification, which is difficult to set in practice.
The flexibility of the RVNP error model also gives rise to certain limitations. In RVNP, we are required to choose an error model architecture together with network hyperparameters. Although the neural-network parametrised error model performs more favourably than the global covariance in most of the tasks, there are two issues which can arise from its implementation. Firstly, we fixed the architecture capacity, implying that the neural-network covariance matrix is massively over-parametrised in the $N_{\rm obs}=1$ and $10$ cases. Secondly, as highlighted in the spectra task, the highly nonlinear bridging of the simulation-to-reality through covariance inflation gap may not be suitable if many of the points lie within distribution in a manner that cannot be distinguished by $\bm{\theta}$. In the spectra task, significant subsets of $\Theta$ are mapped to observed points in a manner that the covariance matrix collapsed to a delta on the subset. We believe that a reliable calibration set that covers such subsets can address these collapses. The impact of a complex bridging of the simulation-to-reality gap has also been witnessed in unsupervised domain adaptation (elsemüller2025doesunsuperviseddomainadaptation).
Introducing an error model inherently adds noise into the forward model. We found that the lowest the covariance would decrease to in the well-specified tasks amounted to a standard deviation of $0.01$ on the Gaussian uncertainty in each direction. While small, if the simulator is very sensitive to changes in the underlying parameters this could be consequential. 2021arXiv211006581H observe that SBI methods tend to produce overconfident posterior approximations in general, and perhaps introducing an error model in this situation is less problematic.
6.2 Choice of summary statistics
In this paper, we used summary statistics that are hand-crafted and those that are inferred using a neural statistic estimator (NSE).
Interpretable, hand-crafted summary statistics, or preparing the data into a manageable form with physically motivated units is often highly desirable (kelly2025simulation). In the context of misspecification, hand-crafted summary statistics allow RVNP to be used when inferring the misspecification for the purpose of model criticism, providing an understanding of the discrepancies between the simulator model and the data to the practitioner. Often, scientific discoveries are made in physical units by fitting for the discrepancy between the model and the observed data. For example, in astronomy dust extinction is inferred by comparing the global misspecification between the simulator and the observed spectral data. In RVNP, we return both the posterior and the forward model, allowing us to inspect and criticise the model, but also recover fast and robust posterior inference that can readily be applied to large datasets.
Adopting a neural statistic estimator (NSE) to embed the data into a low-dimensional representation can be highly valuable in SBI if the data is complex and high-dimensional (deistler2025simulationbasedinferencepracticalguide). As RVNP is an extension of neural likelihood estimation (NLE) we cannot jointly learn the NSE and the likelihood proxy in an end-to-end fashion (brehmer2020flows). A downside of this is that we have to specify separate hyperparameters for the optimiser when inferring the NSE. However, staged training allows us to choose objectives that guarantee global sufficiency (chen2021neuralapproximatesufficientstatistics), which may be highly desirable for amortised Bayesian inference. As RVNP pre-trains the NSE, we can use the InfoMax objective (hjelm2018learning) and benefit from its theoretical guarantees of global sufficiency. A staged training allows us to decouple the inference from learning the NSE, and can help the Bayesian workflow if the main computational difficulty is in dealing with the high-dimensional data once simulated. This is true particularly for amortised learning where global sufficient statistics are needed and an up-front simulation budget is defined. Non-amortised methods train in rounds, often making end-to-end methods more appealing.
When learning the NSE and the posterior in an end-to-end fashion, a form of domain adaptation is necessary to account for misspecification. While this will help against the observed data having a very poor representation under the NSE, there is an inherit trade-off between learning the posterior and accounting for the misspecification in a manner that depends on a user-defined hyperparameter. Although RVNP requires more tunable hyperparameters as a result of the variational inference (and the InfoMax objective, if adopted), it requires no parameters that dictate the strength of the misspecification.
Similarly to error modelling, complex domain adaptation in a nonlinear fashion may have infinitely many solutions that can account for the simulation-to-reality gap in a manner that is highly uninterpretable, and offers very little model criticism. A separate body of work should discuss error modelling using calibration set, and compare with the results of wehenkel2025addressing.
@@ 6.3 Prior misspecification
Prior misspecification was not explicitly addressed within the RVNP tasks. Assuming that the neural likelihood proxy is learnt exactly, RVNP should have the same limitations as likelihood-based prior misspecification if the support of the adjusted prior is a subset of the support of the original prior used to train the likelihood. In reality, the ability to train the likelihood is influenced by the prior over $\bm{\theta}$, and it may struggle to learn the distribution in regions where there are a low density of points.
In the experiments, we chose relatively uninformed priors. However, we stress that the pendulum task benefits from the uniform bounds in $\bm{\theta}$ defined to generate the simulator samples. In the pendulum task, the points appear OOD under the misspecification because its effect makes the points appear as if they have a higher fundamental frequency relative to the summary statistics. However, the information about the misspecification is only available to us because we assume that the prior is well-specified and that there is an upper bound to the fundamental frequency. With expert knowledge this is not an unreasonable assumption. However, we would lose the robustness of the inference if we assumed a wider prior on the fundamental frequency, causing the synthetic DGP to cover the observed data points. Therefore, there is a trade-off between extending the prior to higher fundamental frequencies and robustness.
We note that in the Spectra task we defined a wide, poorly-adopted prior relative to the observed data, which occurred unintentionally due to the selection effects of the real data. However, this did not cause any issues. There are, however, hard boundaries along a certain direction due to the sharp cut off expected from observed metallicities.
6.4 Variational Inference and sample-importance-resampling
RVNP builds on variational methods for solving the inverse problem under a learnt likelihood in SBI (glockler2022variational), and relies on a conditional neural density estimator to approximate the likelihood from a fixed budget of simulations from the simulator model. In this sense, RVNP extends NLE by accounting for misspecification using an error model. We invoke variational inference to overcome the computational expense of using Hamiltonian Monte Carlo 2011hmcm.book..113N or other Markov Chain methods to jointly infer the parameters of interest and the parameters of the error model. The sharing of the parameters over a large model makes the inference highly dependent on the gradients, and struggles to converge even for a diagonal Gaussian covariance error model. Furthermore, variational methods allow us to use more expressive models for the error model distribution.
NLE methods require more hyperparamter tuning and model choice by adopting the variational inference posterior, which also comes at an increased computational expense (deistler2025simulationbasedinferencepracticalguide). However, if the amortised likelihood has been trained well, variational methods with sample-importance-resampling can compete with the accuracy of NPE (glockler2022variational). Importantly for the misspecification problem, while we increase the number of tunable parameters and architecture choices for the inference step, we do not adopt hyperparameters that control the strength of the misspecification. In this situation, the difficult arises from the inference step as opposed to accounting for the misspecification. We argue that this is sensible for misspecification.
In general, we find that sample-importance-resampling is important for making the inference more robust. In particular, it is useful for "cleaning samples" that the posterior has generated which are not compatible with the prior distribution or the likelihood. We present an example of this in Figure LABEL:spectra_posterior_comparison_sir.